OpenStudy (anonymous):

Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form. Please help me.

OpenStudy (anonymous):

saifooooooo!

OpenStudy (saifoo.khan):

\[y - y_1 = m(x-x_1)\]

OpenStudy (saifoo.khan):

@satellite73 !! the only 100th user! :D Congrats and good morning! :)

OpenStudy (saifoo.khan):

\[y - 4 = 2(x-(-3))\]Solve.

OpenStudy (anonymous):

really? yesterday i was at 99 i lost 10,000 medals too! how about you?

OpenStudy (anonymous):

im bad at solving equations lol

OpenStudy (anonymous):

y-4=2(x+3)

OpenStudy (anonymous):

\[y-4=2(x+3)\] distributive law \[y-4=2x+6\] add 4 to both sides \[y=2x+10\]

OpenStudy (saifoo.khan):

i have only 4852 now. Someone stole them! :D hehe. ;) No worries, we will earn more. haha. But you are only 100th ranker! :D

OpenStudy (saifoo.khan):

@satellite73